OpenStudy (boothepest1):

Judi is 9 years older than Baz. Baz is twice as old as Caleb. All their ages add up to 64. What are their ages? I'm in highschool we are doing Throwback Thursday in Math. I feel silly but I'm too exhausted to think. XD

OpenStudy (anthonyym):

Let J stand for Judi, B stand for Baz. J = B + 9 B = 2C

OpenStudy (anthonyym):

Oops and also J + B + C = 64

OpenStudy (anthonyym):

Any questions on that part

OpenStudy (boothepest1):

@anthonyym No thanks, I'm good. I'm just exhausted I have to finish my lesson but I'm too tired.

OpenStudy (anthonyym):

alright you got the rest?

OpenStudy (boothepest1):

@anthonyym Yep thanks, I'll medal!